List of Publications

Dmitry V. Khmelev

October 9, 2001

PUBLICATIONS IN REFEREED JOURNALS

1. Khmelev D.V. Reconstruction of linear index expressions for reducing programs to a linear class//Zhurnal Vychislitel’noĭ Matematiki i Matematicheskoĭ Fiziki. Rossiĭskaya Akademiya Nauk, 1998, vol.38, pp.532–544. Translated in “Computational Mathematics and Mathematical Physics”, 1998, vol. 38, No.3, pp.513–525.

2. Khmelev D.V. Limiting theorems for asymmetrical transportation networks//Fundamental and Applied mathematics, 2001, vol.7, no.4, pp. 1259– 1266

3. Oseledets V.I., Khmelev D.V. Stochastic transportation networks and the stability of dynamical systems//Teoriya Veroyatnosteĭ i ee Primeneniya, 2001, v.46, No.1, pp.147–154. Translated in ‘Theory of Probability and its Applications”.

4. Oseledets V.I., Khmelev D.V. Global stability of infinite systems of nonlinear differential equations, and non-homogeneous countable Markov chains (in Russian)[Global’naya ustoichivost’ beskonechnykh system nelineinykh differencial’nykh uravnenii]//Problemy Peredachi Informatsii, 2000, vol.36, no.1, pp.60–76. Translated in “Problems of Information Transmission”, 2000, vol.36, no.1, pp.54–70

5. Khmelev D.V. Markov chains and classification of text passages. (in Russian) [Raspoznavaniye avtora teksta s ispol’zovaniyem cepei A.A. Markova] Vestnik MGU, Ser.9:Philologiya, No 2, 2000, pp. 115–126.

6. Khmelev D.V. Disputed Authorship Resolution through Using Relative Empirical Entropy for Markov Chains of Letters in Human Language Text//Journal of Quantitative Linguistics, 2000, Volume 7, No 3, pp.201– 207.

7. Kukushkina O.V., Polikarpov A.A., Khmelev D.V., Using Literal and Grammatical Statistics for Authorship Attribution (in Russian) [Opredeleniye avtorstva teksta s ispol’zovaniyem bukvennoi i grammaticheskoi informacii]//Problemy Peredachi Informatsii, 2001, vol.37, no.2, pp.96–108. Translated in “Problems of Information Transmission”, pp. 172–184.

8. Khmelev D., Tweedie F. Using Markov Chains for Identification of Writers//Literary and Linguistic Computing, 2001, vol.16, no.4, pp.299–307.

9. Khanin K., Khmelev D., Rybko A., Vladimirov A. Steady solutions for FIFO networks//Moscow Mathematical Journal, vol.1, no.3, pp.407–419.

PROCEEDINGS AND PREPRINTS

1. Afanassieva L.G., Khmelev D.V. Large transportation networks with finite-dimensional state space. Asymptotic approach//Proc. of 7th Vilnus Conference on Math. Stat., 1998.

2. Afanassieva L.G., Khmelev D.V. Large transportation networks with a finite-dimensional state space//Teorija Verojatnosteĭ i ee Primenenija 1999, vol.44, N1, p.157–158. Translated to English in “Theory of Probability and its Applications” 1999, vol.44, N1, p.142–143.

3. Khmelev D.V., Oseledets V.I. Mean-field approximation for stochastic transportation network and stability of dynamical system//Technical Report 445, University of Bremen, June 1999. Preprint.

4. Khmelev D., Spodarev E. Stability of infinite closed Jackson networks//XX International Seminar on Stability Problems for Stochastic Models, Lublin – Naleczow, 5-11 September, 1999, Abstracts, p.85.

5. Khmelev D., Polikarpov A.A. Basic Assumptions about a Sign’s Life Cycle for Mathematical Modelling of Language System Evolution//Proceedings of the fourth conference of the International Quantitative Linguistics Association edited by R.H. Baayen. Prague, August 24-26, 2000.

6. Khmelev D. Disputed Authorship Resolution Using Relative Empirical Entropy For Markov Chain of Letters in a Text//Proceedings of the fourth conference of the International Quantitative Linguistics Association edited by R.H. Baayen. Prague, August 24-26, 2000.

7. Khmelev D.V., Spodarev E. Infinite closed Jackson networks. Technical Report Math/Inf/00/07, Friedrich-Schiller-Universität Jena, April 2000.

8. Khmelev D., Spodarev E. “Non-homogeneous zero-range interaction processes”, resume of the contribution to the XX Seminar on Stability Problems for Stochastic Models, Lublin - Naleczow, september 5-11 1999.

9. Khmelev D.V. Complexity Approach to Disputed Authorship Attribution (In Russian: Slozhnostnoi podkhod k zadache opredeleniya avtorstva)//Proceedings of the International Congress “Russian Language: Contemporaneity and Fates in History”, pp.426-427, Moscow, March 2001.

IN PREPARATION

1. Khanin K., Khmelev D. Homeomorphisms of the circle with the breaktype singularity I. Periodic Rotation Numbers. (in preparation, 2001)

2. Khmelev D. Periodic trajectories for homeomorphisms of the circle with two break-type singularities. (in preparation, 2001)

UNPUBLISHED MANUSCRIPTS

1. Khmelev D.V. On program “Repetitor”//http://zhurnal.lib.ru/h/

hmelew_d_w/